OpenStudy (anonymous):

What is the slope-intercept form of this equation? 6x + 4y = 8

OpenStudy (tkhunny):

Have you considered solving for y?

OpenStudy (anonymous):

That is what slope-intercept form means. It is when y = the equation.

OpenStudy (anonymous):

First you can subtract 6x from both sides This will give you 4y = -6x + 8

OpenStudy (anonymous):

You can then divide 4 from each side y = -6/4 + 8/4 y= -3/2 + 2 Does that make sense?

OpenStudy (anonymous):

Yeeeesss Thanksss .

OpenStudy (tkhunny):

Personally, I would make your life simpler by first dividing by 2. No worries. Unique answers do not care how you find them.

OpenStudy (anonymous):

y= -1.5x +2

OpenStudy (haseeb96):

the formula of slope intercept is this ... y=mx+b and the given equation is this,,, 6x+4y=8 now we will change this equation into slope intercept form 4y=8-6x \[y=\frac{ -6x+ 8 }{ 4 }\] \[y=\frac{ -3x }{ 2 } + 2 \] this is the slope intercept equation